OpenStudy (anonymous):

The seats in a theatre are arranged in parellel rows that form a rectangle region. Th number of seats in each row of the theatre is 16 fewer than the number of rows. How many seats are in each row of the theatre that has 1,161 seats ?

OpenStudy (saifoo.khan):

i did this one before!

OpenStudy (saifoo.khan):

let x be the number of seats in a row x(x-16)=1161 x^2-16x-1161=0 x^2-43x+27x-1161=0 x(x-43)+27(x-43) (x+27)(x-43) so x=43 but the option given is 27 so the answer is 27

OpenStudy (anonymous):

x(x - 16) = 1161 = 27 * 43 = 43(43 - 16) Therefore, there are forty-three rows of twenty-seven chairs.